OpenStudy (anonymous):

please help! medal will be given. Ari’s chickens eat their pellets at a constant rate throughout each day. What is the slope of the line that represents the ounces of pellets eaten?

OpenStudy (anonymous):



OpenStudy (anonymous):

@ilikemath50 @StudyGurl14 @WinterRaine sorry for the random names .-.

OpenStudy (studygurl14):

find the slope using two points. do you know the fomrula?

OpenStudy (anonymous):

y=mx+b ? @StudyGurl14

OpenStudy (anonymous):

y=mx+b is correct @helphelpthehippo

OpenStudy (anonymous):

oh okay so the answer is..39? @queencandy

OpenStudy (studygurl14):

no @helphelpthehippo..hold on

OpenStudy (studygurl14):

Slope formula: where \[(x_1,y_1) \] and \[(x_2,y_2)\] are on the same line, the slope equals \[\frac{ y_2-y_1 }{ x_2-x_1 }\]

OpenStudy (anonymous):

okay. @StudyGurl14

OpenStudy (anonymous):

 so.. @StudyGurl14 sorry this is not my best subject.